[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Thu Nov 16, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Thu Nov 16 19:30:45 UTC 2017


HSU Shift Log for Thu Nov 16 2017 14:30:39 GMT-0500 (EST)

Shift: H

Forecaster: M Brennan

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues:

Miscellaneous issues: Submitted a test PSurge run for AL802017 at 15Z.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20171116/81110aa7/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list